OpenStudy (anonymous):

percent of change x to 3x

OpenStudy (lgbasallote):

haha tricky one ;D go @Dockworker :DDD

OpenStudy (anonymous):

\[\frac{3x-x}{x}=\frac{2x}{x}=2=200percent\]

OpenStudy (anonymous):

200% increase

hero (hero):

Oops

OpenStudy (lgbasallote):

haha told you tricky @Hero :P

hero (hero):

Yeah, you're right Dockworker,I was working off the top of my head rather than using the formula. Ill-advised. It wasn't tricky @lgbasallote , just needed to apply the formula

OpenStudy (anonymous):

how do you put the line in between the numerator 3x-x and the denominator x??. :D

OpenStudy (lgbasallote):

haha same thing =)) if it can make someone careless..it's tricky! =)) im wondering bout the equations to..someone said \frac\ thingy

hero (hero):

frac{2x - x}{x} in equation editor

hero (hero):

First set of brackets is numerator, second set of brackets is denominator

OpenStudy (anonymous):

cool, thx :D

hero (hero):

Try it and see what happens